OpenStudy (anonymous):

An investor wants to analyze the earnings of a mutual fund account. Three years ago, the value of the account was $24,000 and it is now worth $31,200. If the account is compared to a bank account paying interest that is compounded continuously, what interest rate rounded to the nearest hundredth of a percent would the bank account have to pay to match the mutual fund account’s earnings? (Assume the only deposit was to open the account.)

OpenStudy (amistre64):

how we doin?

OpenStudy (anonymous):

i need help with this now

OpenStudy (amistre64):

3 years ago, the value of the account was $24,000 and it is now worth $31,200. compounded continuously, A = Pe^(rt) A/P = e^(rt) ln(A/P) = rt ln(A/P) ------ = r t ln(31200/2400) -------------- = rate 3

OpenStudy (amistre64):

i missed a 0 .. spose to e 24000 :)

OpenStudy (amistre64):

what do you get when you plug this into a calculator?

OpenStudy (anonymous):

.262?

OpenStudy (amistre64):

lol ... i get .08745.... round as needed